# Guided simulation of conditioned chemical reaction networks

### v1

@@ 5.1. Simulation of the guided process for underlying processes with time-homogeneous intensities
In many applications, such as the examples in 2.3, the intensities $\lambda_{\ell},\,\ell\in{\mathcal{R}}$ of the underlying process only depend on the state $x$ and don’t have a direct dependence on $t$. Here we consider this case and present a method for simulating the guided process in such a scenario. Note that if the underlying rates are time-dependent, but bounded, these upper bounds can easily be included in the thinning step.
Simulation of a guided process $X\sim\mathbb{P}^{g}$ requires sampling of reaction times with intensities $\lambda^{g}_{\ell}$ given in Equation (3.4). Algorithm 1 is applicable in this case using the bounds from C.1. However, these bounds are quite rough and therefore the acceptance probability is very low, making the algorithm very inefficient.
@@ 5.1.1. Special case: guided process induced by $g_{\varepsilon}$
In case we take $g_{\varepsilon}$, the map $t\mapsto\alpha_{\ell}^{g_{\varepsilon}}(t,x)$ (defined in Equation 4.6) is monotone. Therefore, given $(t,x)$, a tight upper bound can be found at either $(t,x)$ or $(T,x)$. Unfortunately, the upper bound at $T$ is typically far from sharp hampering efficiency in simulation, especially when $\varepsilon$ is chosen to be small. To resolve this issue, we employ Algorithm 2. In lines 22, $\lambda_{\ell}(\cdot,x)$ is increasing in time. However, we can choose $\delta$ such that $t+\delta<T$ and use that, on the interval $[t,t+\delta]$, $\lambda_{\ell}^{g_{\varepsilon}}(\cdot,x)$ is upper bounded by $\lambda_{\ell}^{g_{\varepsilon}}(t+\delta,x)$. Hence, on $[t,t+\delta]$, we apply the thinning property for an inhomogeneous Poisson process with rate $\lambda_{\ell}^{g_{\varepsilon}}(\cdot,x)$ and if no reaction occurs in this interval, we move $t\leftarrow t+\delta$.
@@ 5.1.2. Choice of $\delta$
In Algorithm 2, the choice for $\delta$ in line 2 is arbitrary. In the simulations, we opted to include a minimum acceptance rate. The acceptance rate of a proposed time $\tau_{\ell}$ is given by
@@ 5.2.1. Death model and comparison to Golightly & Sherlock, (2019)
We consider the pure death process of 2.3 and estimate the distribution of $X(T)$. For convenience, we denote by $g^{v}$ a guiding term $g$ chosen for the conditioning $X(T)=v$. By 3.5, for any $v\in{\mathbb{S}}$
$p(0,x_{0};T,v)=g^{v}(0,x_{0}){\mathbb{E}}^{g^{v}}\left[\frac{\Psi_{T}(X)}{g^{v}(T,X(T-))}\mathbf{1}_{\{X(T)=v\}}\right]$
@@
$\hat{p}(v)=\frac{1}{N}\sum_{i=1}^{N}g^{v}(0,x_{0})\frac{\Psi_{T}(X_{i})}{g^{v}(T,X_{i}(T-))}\mathbf{1}_{\{X_{i}(T)=v\}}$ (5.4)
It is a well-known result that $X(T)\sim\mathrm{Binom}\left(x_{0},e^{-cT}\right)$, so we will use this density for comparison.In the following experiment, we use (5.4) to estimate the density of the $\mathrm{Binom}\left(x_{0},e^{-cT}\right)$-distribution with $x_{0}=50$, $T=1$ and $c=1/2$ with the functions $g=g_{\mathrm{LNA}}$ from the LNA method with restart, see (4.1), $g=g_{\varepsilon}$ chosen using a scaled diffusion, see (4.4), and $g$ chosen using the density of a unit rate poisson process with intensity $\theta$. Given, $v$, the value for $a=\sigma^{2}$ (appearing in $g_{\varepsilon}$) is chosen around $2.5(x_{0}-v)$, which was rougly the optimum of the map $a\mapsto g^{v}(0,x_{0})\Psi_{T}(X)\mathbf{1}_{\{X(T)=v\}}/g^{v}(T,X(T-))$ for the $1\%$, $50\%$ and $99\%$-quantiles of $X(T)$ (see Remark 4.7). For $\theta$, we applied Theorem 4.2 of Corstanje et al., (2023) which proves equivalence of the measures when $\theta\leq\min\{cx:x\in\{x_{0},x_{0}-1,\dots,v\}\}=cv$ and chose $\theta=cv$; the largest allowed value. For both the diffusion and the LNA guiding term, we used $C=\varepsilon=10^{-5}$.
@@
For forward simulation under ${\mathbb{P}}^{g}$ we employ Algorithm 2 where, for the diffusion, $\delta$ is chosen according to (5.3). For the LNA and poisson $g$, such a closed form expression cannot be derived, and we thus, given $t<T$, set $\delta=\frac{T-t}{2}$. This does not affect the validity of the algorithm, only the acceptance probability of a sampled reaction time. The estimated probability mass functions as well as the percentages of paths with $X(T)=v$ for each $v$ using (5.4) with $N=15000$ are plotted in Figure 3. The sum of squared errors $\sum_{k}(q(v)-\hat{p}_{v})^{2}$, with $q$ denoting the probability mass function of the $\mathrm{Binom}(x_{0},e^{-cT})$-distribution, for each method is shown in table Table 1.
Figure 3. Left: Estimates of the probability mass function of $X(T)\mid X(0)=x_{0}$ using the guiding functions $f$ from the LNA method, a scaled diffusion and a Poisson process. The upper barplot is the true density. For each $v$, we estimated using (5.4) with $N=15000$. Right: The percentages of sample paths with $X(T)=v$ for each $v$ for the three methods.
Method Squared error
LNA (with restart) $1.6\cdot 10^{-4}$
Diffusion guiding term $4.6\cdot 10^{-4}$
Poisson guiding term $1.1\cdot 10^{-5}$
Table 1. sum of squared errors of the estimates from Figure 3 for each of the methods
@@
From Figure 3 and Table 1, we first observe that the panel on the left hand side of Figure 3 appears to confirm $\mathbb{P}_{T}^{h}\ll\mathbb{P}_{T}^{g}$ for all choices of $g$. Moreover, we see that the overall error of the Poisson guiding term is lowest and the diffusion guiding term performs particularly well for low values of $v$ for which the process has to make a lot of jumps. However when conditioning on high values, the amount of sample paths that end up in the right point is low, leading to a larger variance in the estimator. The performance of the Poisson guiding term is explained due to the typically lower guiding term, which is 00 when the conditioning is reached while the other two guiding terms can go past it. Since the measures are equivalent for the Poisson, all processes end up in $v$.
@@ 5.2.2. Gene transcription and translation
We start with the GTT-model presented in 2.4. Figure 4 contains realizations of the process $X$ under ${\mathbb{P}}^{g}$ conditioned to hit $(1,11,56)$ at time $T=1$. We used guiding induced by $g_{\varepsilon}$ with $\varepsilon=10^{-5}$ and $a=a_{\mathrm{CLE}}(0,x_{0})$. The plot contains a random sample of 10 out of 1000 simulated trajectories. All trajectories satisfy $X(T)=(1,11,56)$.
Figure 4. Realization of a guided process starting from $x_{0}=(1,50,10)$ conditioned to be at $(1,10,50)$ at time $T=1$ of the GTT-model from 2.4 with $\kappa_{1}=100$, $\kappa_{2}=10$, $d_{M}=25$ and $d_{P}=1$. The thick line is the original (unconditioned) process.
For Figure 5, we sampled 15 times at random and saved the values of randomly chosen components of a forward simulation to simulate a guided process through multiple partial observations. All of the 1000 trajectories hit all partial observations.
@@ 5.2.3. Enzyme kinetics
As described in 2.5, this is a very interesting example as it contains a monotone component as well as an absorbing state, where the process dies. We use this example to compare various guiding terms. We choose the starting point $x_{0}=(12,10,10,10)$ and parameters $(\kappa_{1},\kappa_{2},\kappa_{3})=(5,5,3)$. We consider three scenarios in which the LNA method with restart, the guiding term obtained from a scaled diffusion presented in Section 4 and the guiding term in which one of the components is replaced by a Poisson guiding term as presented in Section 4.4.
@@
First we simulate 100 trajectories of each process and check the percentage of paths that satisfy $X(T)=x_{T,q}$ for $q=0.01,0.5,0.99$. We used the same $\delta$’s as with the death process. For the diffusion guiding term, we took $a=a_{\mathrm{CLE}}(0,x_{0})$. For the Poisson guiding term, we took for the non-monotone components $3\times 3$ with rows $1-3$ and columns $1-3$ of $a_{\mathrm{CLE}}(0,x_{0})$ for $a$ and $\lambda_{3}(0,x_{0})$; the lower bound of the reaction intensity for the reaction that induces the monotone component for the intensity of the Poisson component. The results are summarized in Figure 6.
@@
Just as with the death process in Figure 3, we see that the diffusion guiding term and the LNA guiding term struggle when the monotone component is conditioned not to make many jumps.
@@ 6. Discussion
In this paper we have provided sufficient conditions for constructing valid guided processes, where “valid” refers to $\mathbb{P}^{h}_{T}\ll\mathbb{P}^{g}_{T}$. We have presented various choices of $g$ and conclude that among those there is no best choices in terms of closeness to true conditioned processes and computational costs. When used within a sequential Monte Carlo or Markov chain Monte Carlo algorithm, a mixture of proposals may be beneficial. The inherent discreteness of chemical reaction processes makes it a hard problem, but it works to our advantage in the sense that guided processes can be constructed on $[0,T]$ where $g(T,x)$ is well defined. This is accomplished by using a guiding term derived from conditioning a diffusion process that is observed with (small) extrinsic noise. The guiding term $g_{\varepsilon}$ has the additional advantage that monotonicity can be exploited for simulating paths ${\mathbb{P}}^{g}$ efficiently.
In the analogous problem of continuous-discrete smoothing for diffusion processes (see e.g. Mider et al., (2021), Beskos et al., (2008), Golightly & Wilkinson, (2008)) there exists a simple random-walk like MCMC-sampler on path space to update guided processes: the preconditioned Crank-Nicolson scheme. Unfortunately, we are not aware of a similar construct for chemical reaction processes or any kernel that leaves the smoothing distribution invariant.
@@ 7. Acknowledgments
We thank M. Schauer (Gothenburg University and Chalmers University of Technology, Sweden) for stimulating discussions on the simulation algorithms.
@@ 8. Declarations
Appendix B Proofs for Section 4.3
@@
    For all fixed $j$, $\Psi_{t}(X)\mathbf{1}_{A_{j}[X^{t}]}$ is ${\mathbb{P}}^{g}$-almost surely uniformly bounded in $t$.
@@
${\mathbb{E}}\left[f(X)\frac{{\mathbb{E}}\left[g(t_{k},X(t_{k}))\mathbf{1}\{L_{k}X(t_{k})=v_{k}\}\right]}{h(0,x_{0})}\mathbf{1}_{A[X^{t_{k}}]}\right]={\mathbb{E}}^{g}\left[f(X)\frac{g(0,x_{0})}{h(0,x_{0})}\Psi_{t_{k}}(X)\mathbf{1}_{A[X^{t_{k}}]}\right]$ (B.1)
@@
${\mathbb{E}}^{h}\left[f(X)\mathbf{1}_{A[X^{t_{n}}]}\right]={\mathbb{E}}^{g}\left[f(X)\frac{g(0,x_{0})}{h(0,x_{0})}\Psi_{t_{n}}(X)\mathbf{1}_{A[X^{t_{n}}]}\right].$ (B.2)
@@
Then (B.1a) is satisfied by construction. Lemmas B.2-B.4 prove the remaining conditions. ∎
@@
Upon defining $Z(t)=g(t,X(t))$, it follows from dominated convergence and 4.10 that for any bounded continuous function $f$,
@@
Note that in (B.6), the convergence/divergence is at an exponential rate and therefore, by (B.6) and 4.14,
@@
for all $s<t_{k}$ whenever $L_{k}x\neq v_{k}$. Hence, if $L_{k}X(s)\neq v_{k}$ for any $s<t_{k}$, $X$ jumps before time $t_{k}$ with probability $1$. Since this holds for any $s<t_{k}$, $X$ jumps an infinite amount of times if $L_{k}X$ does not hit $v_{k}$. We now consider the jumps that take $L_{k}X$ closer to $v_{k}$ with respect to the metric $d_{k}$. Let
@@
By 4.14, there are nonzero entries in this summation that tend to $\infty$ exponentially as $u\uparrow t_{k}$ and thus

### v2

@@ 6.1. Simulation of the guided process for underlying processes with time-homogeneous intensities
In many applications, such as the examples in 2.2, the intensities $\lambda_{\ell},\,\ell\in{\mathcal{R}}$ of the underlying process only depend on the state $x$ and don’t have a direct dependence on $t$. Here we consider this case and present a method for simulating the guided process in such a scenario. Note that if the underlying rates are time-dependent, but bounded, these upper bounds can easily be included in the thinning step.
@@ 6.1.1. Special case: guided process induced by $g_{\varepsilon}$
We consider the guided process with $\lambda_{\ell}^{g_{\varepsilon}}=\alpha^{g_{\varepsilon}}_{\ell}\lambda_{\ell}$, where $\alpha^{g_{\varepsilon}}_{\ell}$ is defined in Equation 4.8. As the map $t\mapsto\alpha_{\ell}^{g_{\varepsilon}}(t,x)$ is monotone, given $(t,x)$, an upper bound can be found at either $(t,x)$ or $(T,x)$. Unfortunately, the upper bound at $T$ is typically far from sharp, hampering efficiency, especially when $\varepsilon$ is chosen to be small. To resolve this issue, we employ Algorithm 2. In lines 22, $\lambda_{\ell}^{g_{\varepsilon}}(\cdot,x)$ is increasing in time. Thus we can choose $\delta$ such that $t+\delta<T$ and use that, on the interval $[t,t+\delta]$, $\lambda_{\ell}^{g_{\varepsilon}}(\cdot,x)$ is upper bounded by $\lambda_{\ell}^{g_{\varepsilon}}(t+\delta,x)$. Hence, on $[t,t+\delta]$, we apply the thinning property for an inhomogeneous Poisson process with rate $\lambda_{\ell}^{g_{\varepsilon}}(\cdot,x)$ and, if no reaction occurs in this interval, we move $t$ to $t+\delta$. Typically, once $LX_{t}=v$ for $t$ near $T$ the map $s\mapsto g(s,x+\xi_{\ell})/g(s,x)$ will be decreasing for choices of $g$ considered in this article. If at line 2 $LX_{t}\neq v$ and $T-(t+\delta)\leq\varepsilon$ we reject the sampled path for specified small $\varepsilon>0$.
@@ 6.1.2. Choice of $\delta$
We now explain that $\delta$ can be chosen such that a minimum acceptance rate is attained.
The acceptance rate of a proposed time $\tau_{\ell}$ is given by
@@ 6.2.1. Death model and comparison to Golightly & Sherlock, (2019)
We consider the pure death process of 2.2 and estimate the distribution of $X(T)$ conditional on $X(0)=x_{0}$. For convenience, we denote by $g^{v}$ the guiding term $g$ chosen for the conditioning $X(T)=v$. By 3.6, for any $v\in{\mathbb{S}}$
$p(0,x_{0};T,v)=g^{v}(0,x_{0}){\mathbb{E}}^{g^{v}}\left[\frac{\Psi_{T}^{g^{v}}(X)}{g^{v}(T,X(T))}\mathbf{1}_{\{X(T)=v\}}\right]$
@@
$\hat{p}(v)=\frac{1}{N}\sum_{i=1}^{N}g^{v}(0,x_{0})\frac{\Psi_{T}^{g^{v}}(X_{i})}{g^{v}(T,X_{i}(T))}\mathbf{1}_{\{X_{i}(T)=v\}}$ (6.4)
It is a well-known result that $X(T)\sim\mathrm{Binom}\left(x_{0},e^{-cT}\right)$, so we will use this density for comparison. In the following experiment, we use (6.4) to estimate the mass function of the $\mathrm{Binom}\left(x_{0},e^{-cT}\right)$-distribution with $x_{0}=50$, $T=1$ and $c=1/2$. We consider four choices for $g$:
  • $g_{F}$ as in (4.1) with $C=10^{-5}$ for $v<32$ and $C=0.3$ for $v\geq 32$ (we comment on this choice of $C$ below).
  • $g_{\mathrm{LNA}}$ from the LNA method with restart as in (4.2), with $C=10^{-5}$.
  • $g_{\varepsilon}$ chosen using a scaled diffusion as in (4.4). We took $\varepsilon=10^{-5}$. For a given value of $v$, the tuning parameter $a=\sigma^{2}$ appearing in $g_{\varepsilon}$ was chosen as $2.5(x_{0}-v)$. This was found to be roughly the optimum of the map $a\mapsto g^{v}(0,x_{0}){\mathbb{E}}^{g_{\varepsilon}}\Psi_{T}^{g^{v}}(X)\mathbf{1}_{\{X(T)=v\}}/g^{v}(T,X(T))$, where the expectation was estimated using $100$ forward simulated paths and $v$ was taken to be the $1\%$, $50\%$ and $99\%$-quantiles of $X(T)$.
  • $g$ chosen using the density of a reversed (decreasing) constant rate Poisson process with intensity $\theta$. We took $\theta=cv$ (this choice ensures equivalence by Theorem 4.2 of Corstanje et al., (2023)).
@@
For forward simulation under ${\mathbb{P}}^{g}$ we employed Algorithm 2 where for $g_{\varepsilon}$, $\delta$ was chosen according to (6.3). For the other choices of $g$ such a closed form expression cannot be derived, and for $t<T$ we simply took $\delta=\frac{T-t}{2}$. This does not affect the validity of the algorithm, only the acceptance probability of a sampled reaction time.
Results: We took Monte-Carlo sample size $N=15000$. The estimated probability mass functions are depicted in Figure 3. This figure confirms that $\mathbb{P}_{T}^{h}\ll\mathbb{P}_{T}^{g}$ for all choices of $g$ considered. The percentages of paths conditioned on $X(T)=v$ and actually ending in $v$ are depicted in Figure 4 for a range of values of $v$. We now comment on the choice of $C$ for the Fearnhead guiding term. For $C=10^{-5}$ we noticed that conditioning on high values of $v$ caused the guided process to jump very closely to $T$ with high probability. This resulted in numerical instability when computing the integral appearing in $\Psi_{T}^{g}$. For that reason, we took a larger value of $C$ in case $v$ is large. A side effect of that is that there is less strong forcing to hit $v$ at time $T$ which explains the lower fractions of paths ending in $v$ for $v\geq 32$.
@@
We observe that the overall error of the Poisson guiding term is lowest and the diffusion guiding term performs particularly well for low values of $v$ for which the process has to make a lot of jumps. However, when conditioning on high values, the amount of sample paths that end up in the correct state is low, leading to a larger variance of $\hat{p}(v)$ defined in Equation 6.4. The performance of the Poisson guiding term is explained due to the typically lower guiding term, which is exactly 00 when the conditioning is reached at a time prior to $T$. This property is not shared by the other choices considered.
In Table 1 we report the mean squared errors $\frac{1}{\#v}\sum_{v}(q(v)-\hat{p}(v))^{2}$, with $q$ denoting the probability mass function of the $\mathrm{Binom}(x_{0},e^{-cT})$-distribution.
Figure 3. Estimates of the probability mass function of $X(T)\mid X(0)=x_{0}$ using the guiding functions $f$ from the LNA method, a scaled diffusion and a Poisson process. The upper barplot is the true density. For each $v$, we estimated using (6.4) with $N=15000$.
Figure 4. The percentages of sample paths with $X(T)=v$ for each $v$ for the three methods with the same settings as Figure 3
Table 1. Mean squared errors of the estimates from Figure 3 for each of the methods
Method Mean Squared Error
Fearnhead $1.5\cdot 10^{-5}$
LNA (with restart) $6.5\cdot 10^{-6}$
Diffusion guiding term $2.1\cdot 10^{-5}$
Poisson guiding term $4.0\cdot 10^{-7}$
@@ 6.2.2. Gene transcription and translation
We consider the GTT-model presented in 2.3. Figure 5 contains realizations of the process $X$ under ${\mathbb{P}}^{g}$ conditioned to hit $(1,11,56)$ at time $T=1$. We used guiding induced by $g_{\varepsilon}$ with $\varepsilon=10^{-5}$ and $a=a_{\mathrm{CLE}}(0,x_{0})$. The plot shows 10 sampled trajectories. Out of $1000$ more trajectories sampled, all of these end in the correct point $X(T)=(1,11,56)$.
Figure 5. Realization of a guided process starting from $x_{0}=(1,50,10)$ conditioned to be at $(1,10,50)$ at time $T=1$ of the GTT-model from 2.3 with $\kappa_{1}=100$, $\kappa_{2}=10$, $d_{M}=25$ and $d_{P}=1$. The thick line is the original (unconditioned) process.
Next, from a simulated forward path we chose 15 times at random and saved the values of randomly chosen components of the process. Taking these values as observations we show in Figure 6 multiple realisations of the guided process. Out of $1000$ simulated trajectories we found that all of those passed through all partial observations.
@@ 6.2.3. Enzyme kinetics
As described in 2.4, this is a very interesting example as it contains a monotone component as well as absorbing states, where no more reactions can occur. For example, the state $X(t)=(0,20,0,32)$ can be reached with a sequence of reactions from $x_{0}$, but no further reactions are possible in this state. We use this example to compare various guiding terms. We choose the starting point $x_{0}=(12,10,10,10)$ and parameters $(\kappa_{1},\kappa_{2},\kappa_{3})=(5,5,3)$. We consider three scenarios in which the LNA method with restart, the guiding term obtained from a scaled diffusion presented in Section 4 and the guiding term in which one of the components is replaced by a Poisson guiding term as presented in Section 4.3.
@@
First we simulate 100 trajectories of each process and check the percentage of paths that satisfy $X(T)=x_{T,q}$ for $q=0.01,0.5,0.99$. We used the same scheme for simulating the guided process as for the death process. For the diffusion guiding term, we took $a=a_{\mathrm{CLE}}(0,x_{0})$. For the Poisson guiding term, the auxiliary process from which $g$ is obtained contains a scaled diffusion in the $3$ components and a Poisson processes in the third component. The diffusion is scaled by a matrix $a$, in which we used the first three rows and the first three columns of $a_{\mathrm{CLE}}(0,x_{0})$. The intensity of the Poisson component was taken as $\lambda_{3}(0,x_{0})$, which is a lower bound of the reaction intensity for the reaction that induces the monotone component. The results are summarised in Figure 7.
@@
Similar to the death process in Figure 3, we see that the diffusion guiding term and the LNA guiding term struggle when the monotone component is conditioned not to make many jumps.
We used the same method as for the death processes to estimate $p(0,x_{0};T,v)$ for scenario’s A, B and C. In each scenario, we computed $\hat{p}$ through (6.4). In each scenario, we make 200 estimates for $p(0,x_{0};T,v)$ by computing $\hat{p}$ with $N=1000$ processes and we compared the LNA method (without restart cf. Section 4.3 of Golightly & Sherlock, (2019)) with the diffusion guiding term with $\varepsilon=10^{-5}$ and $a=100a_{\mathrm{CLE}}(T,x_{T,q})$ for $q=0.01,0.5,0.99$. For the LNA method we assumed extrinsic noise $C=2000I$, $C=500I$ and $C=200I$ for scenario’s A, B and C, respectively. Histograms of the estimates are given in Figure 8, Figure 9 and Figure 10. The MSEs of the estimates are summarised in Table 2.
Figure 8. Histograms of $200$ estimates for $p(0,x_{0};T,v)$ in scenario A using the LNA method without restart (top) and and the diffusion method (bottom). The true value was estimated using 10,000 forward simulations.
Figure 9. Histograms of $200$ estimates for $p(0,x_{0};T,v)$ in scenario B using the LNA method without restart (top) and and the diffusion method (bottom). The true value was estimated using 10,000 forward simulations.
Figure 10. Histograms of $200$ estimates for $p(0,x_{0};T,v)$ in scenario C using the LNA method without restart (top) and and the diffusion method (bottom). The true value was estimated using 10,000 forward simulations.
Table 2. Mean Squared Error (MSE) of the estimates from Figures 8, 9 and 10 for each of the methods
Method MSE scenario A MSE scenario B MSE scenario C
LNA (without restart) $8.8\cdot 10^{-6}$ $2.3\cdot 10^{-3}$ $1.3\cdot 10^{-3}$
Diffusion guiding term $9.0\cdot 10^{-6}$ $1.8\cdot 10^{-3}$ $1.6\cdot 10^{-3}$
@@ 7. Discussion
In this paper we have provided sufficient conditions for constructing valid guided processes, where “valid” refers to the law of the true conditioned process being absolutely continuous with respect to the law of the guided process induced by $g$ with the laws as defined in Section 3. We have presented various choices of $g$ and conclude that among those there is no best choice in terms of closeness to the true conditioned process and computational cost. When used within a sequential Monte Carlo or Markov chain Monte Carlo algorithm, a mixture of proposals may be beneficial. The inherent discreteness of chemical reaction processes makes it a hard problem, but it works to our advantage in the sense that guided processes can be constructed on $[0,T]$ where $g(T,x)$ is well defined. This is accomplished by using a guiding term derived from conditioning a diffusion process that is observed with (small) extrinsic noise. The guiding term defined in Section 4.3 has the additional advantage that monotonicity can be exploited for simulating paths of the guided process efficiently.
In the analogous problem of continuous-discrete smoothing for diffusion processes (see e.g. Mider et al., (2021), Beskos et al., (2008), Golightly & Wilkinson, (2008)) there exists a simple random-walk like MCMC-sampler on path space to update guided processes: the preconditioned Crank-Nicolson scheme. Unfortunately, we are not aware of a similar construct for chemical reaction processes.
@@ 8. Acknowledgments
The authors would like to thank the anonymous referees for their helpful comments. We also thank M. Schauer (Gothenburg University and Chalmers University of Technology, Sweden) for stimulating discussions on the simulation algorithms.
@@ 9. Declarations
Appendix B Proofs for Section 5
@@
    For all fixed $j$, $\Psi_{t}^{g}(X)\mathbf{1}_{A_{j}[X^{t}]}$ is ${\mathbb{P}}^{g}$-almost surely uniformly bounded in $t$.
@@
${\mathbb{E}}\left[f(X)\frac{{\mathbb{E}}\left[g(t_{k},X(t_{k}))\mathbf{1}\{L_{k}X(t_{k})=v_{k}\}\right]}{h(0,x_{0})}\mathbf{1}_{A[X^{t_{k}}]}\right]={\mathbb{E}}^{g}\left[f(X)\frac{g(0,x_{0})}{h(0,x_{0})}\Psi_{t_{k}}^{g}(X)\mathbf{1}_{A[X^{t_{k}}]}\right]$ (B.1)
@@
${\mathbb{E}}^{h}\left[f(X)\mathbf{1}_{A[X^{t_{n}}]}\right]={\mathbb{E}}^{g}\left[f(X)\frac{g(0,x_{0})}{h(0,x_{0})}\Psi_{t_{n}}^{g}(X)\mathbf{1}_{A[X^{t_{n}}]}\right].$ (B.2)
@@
Then (B.1a) is satisfied by construction. Lemmas B.3-B.5 prove the remaining conditions. ∎
Lemma B.2.
$\lim_{t\uparrow t_{k}}g(t,x)=\begin{cases}g(t_{k},x)&\text{if }L_{k}x=v_{k}\\ 0&\text{otherwise}\end{cases}.$ (B.4)
Proof.
$(v(t)-L(t)x)^{\prime}M(t)(v(t)-L(t)x)=\frac{d_{k}(v_{k},L_{k}x)^{2}}{(t_{k}-t)}+o(t_{k}-t),$ (B.5)
where $d_{k}$ is the metric on ${\mathbb{R}}^{m_{k}}$ defined as
$d_{k}(x,y)=\sqrt{(y-x)^{\prime}\left(L_{k}a_{k}L_{k}^{\prime}\right)^{-1}(y-x)},\qquad x,y\in{\mathbb{R}}^{m_{k}}.$ (B.6)
which yields (B.4). Since $v$ and $L$ are piecewise constant and $\frac{\,\mathrm{d}M}{\,\mathrm{d}t}=M(t)\begin{pmatrix}L_{i}a_{k}L_{j}^{\prime}\end{pmatrix}_{i,j=k}^{n}M(t)$ is positive semidefinite, $\partial_{t}\log g(t,x)\leq 0$. ∎
@@
Upon defining $Z(t)=g(t,X(t))$, it follows from dominated convergence and B.2 that for any bounded continuous function $f$,
@@
Note that in (B.9), the convergence/divergence is at an exponential rate and therefore, by (B.9) and 5.4,
@@
for all $s<t_{k}$ whenever $L_{k}x\neq v_{k}$. Hence, if $L_{k}X(s)\neq v_{k}$ for any $s<t_{k}$, $X$ jumps before time $t_{k}$ with probability $1$. Since this holds for any $s<t_{k}$, $X$ jumps an infinite amount of times if $L_{k}X$ does not hit $v_{k}$.
We now consider the jumps that take $L_{k}X$ closer to $v_{k}$ with respect to the metric $d_{k}$. Let
@@
By 5.4, there are nonzero entries in this summation that tend to $\infty$ exponentially as $u\uparrow t_{k}$ and thus